Operations Report – November 11th

Crew 182 Operations Report  11Nov2017

SOL: 8

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: It is not working.  In the morning, we had to take a shower with cold water. The water heater is not heating the water.  We opened the entire water pipe, but it still did not heat.

COMMUNICATION RADIOS: The constantly noise of one of the radios stops.
SECONDARY AIRLOCK: The door of the secondary airlock is gotten out-off-square.  I am saying that it is wonky since we arrived on Mars.
ATV’s: The blue ATV is not working. When we try to turn it on, it does not answer.

Generator—  It worked the whole night since 6pm of 10Nov2017 until 8:00 am of 11Nov2017.  Turn on at 5.15 pm 11Nov2017.  Operand.

Solar—

READING FROM 11 NOV 2017

 

TIME SOC READINGS
9:11 am 100%
10:20 am 100%
11:20 am 100%
12:20 am 100%
13:20 am 100%
14:20 pm 100%
15:50 pm 98%
16:20 pm 97%
16:45 pm 94%
17:04 pm 91%

 

Diesel – 55 %

Propane – 50 psi

Gasoline (5 Gallon containers for ATV) – 1/4 gallons. In addition the tank of gasoline inside the ATV’s is filled.

Water (trailer) –  0 gallons

Water (static) – 258  gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  36.6 gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes

HabCar used? No

Notes and Comments:
Today we had an advance in the topic about recycling water. The number of gallons of water that we used today is less than the gallons of water used yesterday.  And, I know that tomorrow is going to be better.
Summary of  Hab operations:

We receive a new digester for the toilet.  And we wait that it removes the bad smell of the toilet. We will keep you post.

The area of the tools in the Hab was a mess. Today we clean and organize this place.  Also, in the afternoon, we flushed the toilet with the new digester.   From another hand, we are going to receive the communication radios on Monday.
Finally, today we received some food.  This is great for us. Thanks¡

Summary of GreenHab operations:

We turned on the heater at 18:12 pm with a temperature of 47 °F.  At the same time, we are cleaning this area to make it beautiful.

Summary of ScienceDome operations:

We received a new shop vac. It has been installed in the laboratory. Also, there were cables going through the laboratory that we order to keep the security in the laboratory.

Summary of health and safety issues:

The HOS officer needs two fire alarms and lamps for the science dome and for the greenhouse. In the morning we fixed the stair of the secondary airlock.  Now, that stair is more stable.

Questions, concerns and requests to Mission Support:

We need you to repair the heater water, since it still does not work.

For the next week at the most on Monday, we are going to need water.

I have a doubt, which is the minimum limit for the propane tank and the diesel tank?

Where does the used water go?

When are we going to receive water?

GreenHab – November 11th

Sol 8 – 11/11/2017

Green Hab Officer Atila Meszaros

Ambient: heating 45 F , turned on at 18:12

Ambient with window/door open: door opened at 08:55; closed at 17:39

Shade cloth on/off: on

Average temperature:

Time 09:00 13:00 17:05 18:42
Temperature inside (°C) 26 34 22 13
Temperature outside (°C) 3 6 11 1
% Humidity 16 16 16 16

 

Changes to crops:  Second day of germination! Crops keep growing!

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 09:00 and 13:00

Changes to research plants: none.

Daily watering and amount of water used: 1 L approximately

Narrative:

Hi, Gael! Atila in here, like the Hun. I hope you are having a terrific time.

Today was the day after the germination of the crops from a former experiment that is on the Green Hab (GH).  The seeds keep germinating and the seedling keep growing.  I turned on the heater (45 F).

Species (common name) Number of bowls Number of seedlings
Tomatoes 25 8
Cucumber 7 1
Melon 8 7
Beans 17 7
Pepper 9 3

Tomorrow as soon as possible I’ll update this numbers. It will be probably easier if I make a Drive with the dates and number of emerging seeds. Let me know if you have like a Data Base where I can put this kind of data.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seeds of the experiment I’m using almost a Liter.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus and mustard (as a control).

The analog soil I’m using is becoming clay, and I’m worry that the seeds are not going to be able to break the soil and emerge. So, I’m going to try a new procedure: I’m going to germinate the seeds on Petri dishes and observe growing behavior on those ones.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there!

Have a beautiful night on Earth.

Support/supplies needed: none

Sol Summary – November 11th

MDRS Daily Summary Report for sol 0.8

Summary Title: Reflecting about sustainability.

Mission Status: Our day was concentrated in search to prepare our mission plan, reflects about each individual research plan, and evaluates what kind of meals to be done.

Sol Activity Summary: improving the MDRS.

Look Ahead Plan: Tomorrow, We are planning an EVA walking close to the station.

Anomalies in work:  Nope.

Weather: 10o C, 50o F

Crew Physical Status: Fine.

EVA: No EVAs today.

Reports to be file: We sent in CapComm.

Support Requested: Nope.

Commander Report – November 11th

Yesterday night We had a meeting to clarify some points of our mission and try to improve our results in the next week in MDRS.

6h30m – I woke up.

6h42 – Planning.

6h53m – 8h23m – Walking 90 minutes: 8,6 km or 5,3 miles.

9h10m – Breakfast.

10h20m – Recycling some plastic bottles, and trying to fix something.

12h30m – Lunch: eggs and Hash brown potatoes.

13h – We debated about the mission plan, considering about each research plan. I consider that the main topic of our mission is the sustainability.

15h45m – Camila and Brandon prepared a banana cake.

16h22m 17h02 – Walking 40 minutes: 5,8 km or 3,6 miles. At total I walked today 130 minutes.

Today didn’t happened EVA.

17h20m – I’m working in my research related with sustainability in MDRS:

Sustainability in MDRS:

lessons to Mars research habitats and setlements

Researcher: Prof. Julio Francisco Dantas de Rezende, PhD.

This research searches to answer the question: the Mars Desert Research Station (MDRS) operation can be more sustainable? It is evaluated how environmental, economic, social and personal sustainability issues are presented and is also reflected how MDRS activities would collaborate to Sustainable Development Goals (SDGs), proposing some guidelines to sustainability. It is also important to ask: the results related to Mars would be applied to build a self-sustainable habitats in Earth, mainly in areas affected by climate change as deserts and semiarid regions as can be seen Brazilian Northeast? Reviewing the previous research done at MDRS not was identified any research related to sustainability. Because of that, this research presents a possible great impact to MDRS. It is a challenger identifies the main dimensions that would be considered to evaluates a Mars research station in terms of sustainability and would be considered the great relevance of this research for the future of design of process of future Mars settlements.

The research Sustainability in MDRS: lessons to Mars research habitats and setlements was considered in the Mission Plan prepared today.

19h – CapComm

Operations Report November 10th

Crew 182 Operations Report 10Nov2017

SOL: 7

Name of person filing report: Carmen Atauconcha

Non-nominal systems:

WATER HEATER: It is not working. Today we wanted to take a shower and the water was cold. The heater is not heating the water. We opened the entire water pipe, but it still did not heat.

COMMUNICATION RADIOS: One of the radios is making a special noise constantly. Like every 3-5minutes it sounds, even if we leave it charging.
SECONDARY AIRLOCK: The door of the secondary airlock is gotten out-off-square.
ATV’s: The blue ATV is not working. When we try to turn it on, it does not answer.

Generator— Operating from 09 November, 2017 at 5pm and currently working. 26 hours.

Solar—

READING FROM 10 NOV 2017

TIME SOC READING
8:00 AM – 17:00 PM 100%

READING FROM 09 NOV 2017.

TIME SOC READINGS
8.43 am 98%
9.13 am 98%
9.43 am 98%
10.09 am 100%
10.39 am 100%
12.20 pm 100%
13.01 pm 98%
13.38 pm 96%
13.46 pm 95%
13.56 pm 94%
14.11 pm 93%
14.34 pm 100%
15.19 pm 100%
16.28 pm 98%
17.31 pm 93%

Diesel – 59 %

Propane – 51 psi

Gasoline (5 Gallon containers for ATV) – 1/4 gallons

Water (trailer) – 0 gallons

Water (static) – 275 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 66.8 gallons

Toilet tank emptied: yes

ATV’s Used: RED ATV

Oil Checked: yes

Atv # Fuel Used Gals: 1 gal

Tires Status: Perfect

# Hours the ATVs were Used today: 1 and ½ hours

Deimos used: Yes

Hours: 1 and a ½ hours

Beginning charge: 12:30 pm

Ending charge:

Currently charging: Yes

HabCar used? No

Notes and Comments:
Today we had an advance in the topic about recycling water. The number of gallons of water that we used today is less than the gallons of water used yesterday. And, I know that tomorrow is going to be better.
Summary of Hab operations:

We put the jumper cables delivered on the lower shelf in the EVA room. Also, we fixed the radios because some of them did not have a good sound. In addition, tomorrow we are going to receive new communication radios.

Summary of GreenHab operations:

We are going to heat the greenhouse tomorrow

Summary of ScienceDome operations:

We installed the autoclave perfectly. Even, it has already been used once without any difficulty. At the same time, tomorrow we are going to receive a new shop vac.

Summary of health and safety issues:

The HOS officer needs two fire alarms for the science dome and for the greenhouse.

Questions, concerns and requests to Mission Support:

We need you to repair the heater water, since it still does not work.

For the next week at the most on Monday, we are going to need water.

I have a doubt, which is the minimum limit for the propane tank and the diesel tank?

Where does the used water go?

EVA Report November 10th

EVA Report Sol 7

The EVA started at 10:20 AM instead of 10:00 AM, since I wanted to make sure the crewmembers were properly briefed on radio Comms voice procedure. The red ATV had some trouble starting like yesterday, but once I got the engine running, it didn’t have any problem starting. We drove the rovers to our first location along Cow Dung Road where we took a couple of plant and regolith samples. Afterwards, we proceeded to Tank wash, where we found a large desert shrub that was holding onto a large mound of sand with its root system. EVA operations went smoothly and we obtained all the samples we set out for. Once we returned, we attempted refueling the ATV, but the nozzles on the empty gas containers weren’t compatible with the filled containers, so we couldn’t safely fill up the ATV. Will we have compatible gas nozzles when the ethanol-free gas comes in? Aside from that, we returned early and begun decompression by 12:15 PM.

For missions in hilly terrain, I have to brief the crew members on squatting while walking up or downhill.

Commander Report November 10th

Today is our 7th day. During certain time of my life I had practiced open water swimming. I did some crossing spending more than 3 or 4 hours. Then, my feeling about this mission in MDRS, I would compare as a sea crossing when you are swimming, and also at same time, start to do some statistics in your mind, about how many percent have you swum. In both situations, your mind just to be fine. In our case, in our crossing, the main issue is to keep the good relationship, kindness and respect with each other. Now I’m looking our days here.

Sunday Monday Tuesday Wednesday Thursday Friday Saturday
4
5 6 7 8 9 10 11
12 13 14 15 16 17 18

Something that would be annoying in the space missions is be part of an experiment, or to be compelled to be a subject in a research. This is related when We need to eat some kinds of food or use some kinds of specific products. I can imagine what the astronauts as Scott Kelly, in the International Space Station, would need to deal with it. But, I understand that is all about a “good” cause. If you see, all of us are being a part of a great experiment just agreed by us, with our consent. It is part of a movement to a broader new scientific results.

Yesterday I tried to manage the data space in my computer. It almost full and I need to delete some content and have more space. The mission in MDRS just have had generated a lot of content of photos and videos, hard to process all these data.

In the 7th day I’m just considering MDRS as my home. I like good and bad things here. I just prepared my entire life to this, just traveling to far and urban places, sharing rooms in hostels and visiting different desert areas. I love the landscape in Hanksville.

7h – 8h20m – I did 80 minutes of the exercise. I just walked 8 km or 5 miles. The temperature was -1oC. Watching the second episode of the National Geographic miniserie Mars the crew of Dedalus need to walk in Mars surface after de vehicle broke. The cold would be a high threat to the crew exposed to Extravehicular Activities.

9h40m – Breakfast. Carmen is taking the weight of everyone for her research. Bad news for me: I’m weighing two kilos more. When you travel to space, your body is affected by different ways. So, my diet and exercises here is not helping to loose weight. Good to know just in the middle of the mission and try to adopt new behaviors here. Because of that, is so important keep the control of the different health indicators of an astronaut. I need to be fine with myself and guide the crewmates.

10h25m – The EVA Briefing: Brandon Ferguson, Crew geologist and EVA Lead, proposed some guidelines to communication in ExtraVehicular Activities (EVA). He recommended the use of the Phonetic Alphabet, as can been seen in the next Chart.

Chart 1 – Phonetic Alphabet

A – Alfa H – Hotel O – Oscar V – Victor
B – Bravo I – India P – Papa W – Whiskey
C – Charlie J – Juliet Q – Quebec X – X-Ray
D – Delta K – Kilo R – Romeo Y – Yankee
E – Echo L – Lima S – Sierra Z – Zulu
F – Foxtrot M – Mike T – Tango
G – Golf N – November U – Uniform

Was also proposed some voice procedures to be used in the EVA, as can be seen in the next Chart.

Chart 2 – Voice Procedures in MDRS EVAs

– Affirmative / Affirm = Yes

– Negative = No

– Niner = The Number Nine

– Out = I’ve finished talking to you and do not expect a reply

– Over = I’ve finished talking to you and listening for your reply

– Roger – “I’ve received all of the last transmission”

– Say Again – Repeat

– Wait, over – I don’t have the answer or information on hand. I will attempt to find it, until then, I’ve finished talking and do not expect a reply

– Wilco – will comply (after receiving new directions)

10h40m – EVA with Brandon, Camila and Carmen.

11h – I’m aboard of the Hab. I’m and Atila are supporting Brandon, Camila and Carmen in their EVA. During this time I will try to update my writings. But is no easy. There’s no free time here.

12h – End of the EVA.

14h – Lunch. We recycled beans prepared 2 days ago adding rice and frying everything.

14h30m – I received the information that the interview in

15h – Preparation of commander report.

15h50m – I walked more 40 minutes, 5,7 km or 3,5 miles. At total today, I walked 13,7 km or 8,5 miles.

18h – We debated about how to save water. In the Chart 3 is presented some guidelines.

Chart 3 – Guidelines to Save Water

1 – Communication and Education about save water;

2 – Use body wipes only for shower (unless you have a long hair);

3 – Harvest all dishwashing and cooking water;

Dishwasher => Graywater +> poop

Cookwater => reuse for cooking/redytrating food

4 – Redytrating water:

– reuse for more redytrating or

– grywater

5 – For dishwashing, limit water output to lowest possible setting;

6 – use personal mug;

7 – use of paper towels and wipes to clean up dishes;

8 – Try to let almost possible clean the dish after the meals;

9 – Be quickly when press the flush mechanism;

10 – Don’t shave the beard everyday;

11 – Don’t take a shower everyday;

12 – Not change the clothes everyday.

13 – Storage used water to reuse.

14 – Create recipients to collect water with recycled water bottles, as water cases of 2,5 galons.

15 – Use of hand sanitizer to clean up hand.

I consider the Guidelines to Save Water an important contribution of the mission 182 to MDRS. Our proposition about it is connected with Sustainable Development (SDG) 6: Ensure availability and sustainable management of water and sanitation for all implies. One important topic to MDRS improvements would be invest in a system of water reuse. In International Space Station (ISS) all the water is reused as closed system. MDRS would also like this. An important guideline is have the control about how much water is been used and how much sewage is been generated.

Tomorrow We will be discussing how the activities developed in MDRS would contributes to the implementation of Sustainable Development Goals (SDGs).

The Sustainable Development Goals (SDG) were approved in September 2015 during the Sustainable Development Summit, when it was discussed and approved by the United Nations General Assembly in New York. 193 countries signed the document. The Sustainable Development Goals (SDG) are composed of 17 aims and 169 goals to be achieved by 2030.

Greenhab Report November 10th

Green Hab Officer Atila Meszaros

Sol 7 – 10/11/2017

Ambient (no heating or/cooling): none

Ambient with window/door open: door opened at 10:15; closed at 16:55

Shade cloth on/off: on

Average temperature:

Time 09:50 11:24 12:41 18:42
Temperature inside (°C) 22 31 32 13
Temperature outside (°C) 5 9 14 8
% Humidity 16 16 16 16

Changes to crops: First day of germination!

Daily water usage for crops: 2 galons

Time(s) of watering for crops: 09:50 and 12:41

Changes to research plants: none

Daily watering and amount of water used: 1 L approximately

Narrative:

Hi, Victoria! I’m Atila, like the Hun and I hope you are having a terrific time.

I’m glad to let you know that today was the first day of germination of the crops from a former experiment that is on the Green Hab (GH). I’ll attach you a pic.

Species (common name) Number of bowls Number of seedlings
Tomatoes 25 5
Cucumber 7 1
Melon 8 5
Beans 17 2
Pepper 9 2

Tomorrow as soon as possible I’ll update this numbers. It will be probably easier if I make a Drive with the dates and number of emerging seeds. Let me know if you have like a Data Base where I can put this data. Also, I wanted to ask Shannon how much seeds are in each pot.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seeds of the experiment, I’m using almost a Liter.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus and mustard (as a control).

I think that’s pretty much all if you need more information about the GH I will be glad to send it to you.

Thank you for being there, Vic!

Have a beautiful night on Earth.

Support/supplies needed: none